On January 1, 2021, Maywood Hydraulics leased drilling equipment from Aqua Leasing for a three-year period ending December 31help me with the interest expense, please.
Answer #1

Answer a led Appropriate Entrics for fani, 9021 & DEC 3 1,8213 ->To record the dease, present value do porodic lease paymentsinterest Expense = 289,858 X 6-1, =17392 heaese picapable = 100.000 - 17392 = 84608 Date particulars creditcaal Debit (f) $ 2